# Joint Probability Distribution of mRNA and Protein Molecules in a Stochastic Gene Expression Model

### v2

@@ Abstract
Stochastic modeling of gene expression is a classic problem in theoretical biophysics. However, models formulated via chemical master equation have long been considered analytically intractable unless burst approximation is applied. This article shows that general stochastic gene expression models with an arbitrary number of gene states admit direct analysis. Based on chemical master equation and high-dimensional binomial moment method, we derive recurrence relations for binomial moments in steady state, yielding analytical expressions to arbitrary order in a hierarchical manner. Subsequently, the joint probability mass function of mRNA and protein copy number can be reconstructed. An algorithm is developed for numerical computation. Particularly, explicit expressions for low-order cumulants are presented. Compared with models under burst approximation, the mean remains exact, whereas the variance typically differs. We estimate the difference between two second-order binomial moments using functional analysis, therefore evaluating the validity of burst approximation.
@@ 1 Introduction
Gene expression is one of the most fundamental processes in molecular biology, where the genetic information is realized as functional gene products. In most cases, gene expression consists of transcription followed by translation. During transcription, messenger RNA (mRNA) is synthesized by copying specific segments of DNA, whereas translation is the process of producing proteins using mRNA as templates. To quantitatively understand and predict this process, constructing reasonable mathematical models is necessary. Notably, intrinsic stochasticity is the key feature of gene expression. Such inherent stochasticity originates from random collisions among molecules, and becomes relatively significant due to low numbers of mRNA and protein molecules in most cells [Centraldogmasinglemolecule2011Nature]. Randomized expression of genes can be beneficial for biological adaptation to changing environments [Noisegeneexpression2005Science]. However, how fluctuations propagate within and across cells, and are eventually brought under control, remains an open question.
@@
In general, our primary concern is the mRNA and protein copy number during gene expression, whose dynamics should be modeled as a stochastic process. Note that deterministic approaches like reaction rate equation [AppliedStochasticAnalysis2019] certainly fail because of pronounced stochasticity. More specifically, this stochastic process should be a continuous-time process with a discrete state space. The state space should be discrete because the mRNA and protein counts are nonnegative integers. The simplest class of stochastic processes satisfying both conditions consists of continuous-time Markov chains, whose Kolmogorov forward equation is commonly termed chemical master equation [AppliedStochasticAnalysis2019, StochasticProcessesPhysics2007] in chemical physics literature. In general, chemical master equation is an infinite-dimensional coupled system of ordinary differential equations, and can be equivalently converted into a finite-dimensional partial differential equation system using generating function method or Laplace transform. Therefore, when analyzing a dynamical system described by chemical master equation, neither a universal treatment nor an effective general theory exists, and one must devise techniques tailored to each problem. In particular, compact analytical results are rarely anticipated, and in many cases different kinds of approximation techniques are applied [Approximationinferencemethods2017J.Phys.A-Math.Theor.]. The most widely-used approximation method is diffusion approximation [AppliedStochasticAnalysis2019], which approximates the chemical master equation with a Markov-type stochastic differential equation [BrownianMotionStochastic1998], conventionally termed chemical Langevin equation [chemicalLangevinequation2000J.Chem.Phys.]. This procedure approximates a continuous-time Markov chain with a diffusion process. Therefore, diffusion approximation also becomes inappropriate when the numbers of molecules in the system are extremely low, as the discreteness of the state space becomes significant. Despite the challenges, gene expression models based on chemical master equation are worth in-depth analysis given their biological importance.
@@
Proposing stochastic gene expression models is straightforward. However, it has long been believed that complete gene expression models are analytically intractable. The simplest two-stage model is analyzed in the classic paper [Analyticaldistributionsstochastic2008Proc.Natl.Acad.Sci.U.S.A.]. In the two-stage model, mRNA molecules are continuously produced according to a Poisson process, and translation, hydrolysis of mRNA, and hydrolysis of protein all occur as memoryless, single-step reactions. Surprisingly, in this toy model, analytical distribution of protein copy number is infeasible, since the expression of the generating function contains an incomplete gamma function that cannot be reduced [Analyticaldistributionsstochastic2008Proc.Natl.Acad.Sci.U.S.A.]. Afterwards, very few studies focus on the complete gene expression models. Thereby, existing results about stochastic gene expression models are confined to low-order moments of several toy models [Modelsstochasticgene2005Phys.LifeRev.].
@@
However, an effective approximation technique exists, namely, the burst approximation [Modelsstochasticgene2005Phys.LifeRev., Analyticaldistributionsstochastic2008Proc.Natl.Acad.Sci.U.S.A.]. The burst approximation builds upon the experimental conclusion that mRNAs decay substantially faster than proteins in most cells [Analyticaldistributionsstochastic2008Proc.Natl.Acad.Sci.U.S.A.]. Under burst approximation, the originally gene expression models can be greatly simplified and analytical results become available [Exactdistributionsstochastic2022Phys.Rev.E, Analyticaltimedependentdistributions2023SIAMJ.Appl.Math., Analyticaldistributionsstochastic2008Proc.Natl.Acad.Sci.U.S.A.]. Beyond their roles as approximate substitutes, stochastic gene expression models with translational bursting have their independent theoretical value. Particularly, models with multiple gene states under burst approximation have recently been studied using one-dimensional binomial moment method [Exactdistributionsstochastic2022Phys.Rev.E], and non-Markovian models can also be studied using techniques from queueing theory [Intrinsicnoisestochastic2011Phys.Rev.Lett.]. According to the equivalence between queueing system and gene expression model established in [Solvingstochasticgeneexpression2024Biophys.J.], models under burst approximation can be interpreted as queueing systems with batch arrivals. This perspective may explain the substantial analytical simplifications introduced by the burst approximation, and clarify the theoretical significance of related research. Nevertheless, whether the errors introduced by the burst approximation are controllable has not received much attention. The validity of burst approximation is only evaluated in the simplest two-stage model [Analyticaldistributionsstochastic2008Proc.Natl.Acad.Sci.U.S.A.], and validation for general models also motivates us to gain deeper analysis into complete gene expression models.
@@
Additionally, general numerical methods exist for reaction systems described by chemical master equation. Stochastic simulation algorithm (SSA) [generalmethodnumerically1976J.Comput.Phys.] is a numerical method generating the sample paths of a given continuous-time Markov chain, one realization at a time. Multiple variants of SSA are also available [Stochasticsimulationchemical2007Annu.Rev.Phys.Chem.]. In this article, we employ Python package GillesPy2[GillesPy2biochemicalmodeling2023Lett.Biomath.] to perform SSA. Finite state projection algorithm (FSP) [finitestateprojection2006J.Chem.Phys.] is also a widely-used numerical method, which truncates the chemical master equation and reduces the problem to numerically solving a standard ordinary differential equation system. In general, numerical methods like SSA and FSP are computationally expensive and quickly become infeasible for large systems. Moreover, the aforementioned numerical methods cannot be used to derive further theoretical results. As a result, theoretically analyzing complete stochastic gene expression models is still important.
@@
In this article, we demonstrate that gene expression models can be effectively analyzed without applying burst approximation. Using high-dimensional binomial moment method, we derive recurrence relations for binomial moments of mRNA and protein copy number, yielding exact analytical expressions for binomial moments of arbitrary order. We also design an algorithm that numerically computes, with high accuracy, all binomial moments in a hierarchical way. When binomial moments are obtained, joint probability mass function of mRNA and protein copy number can be constructed directly. In particular, we present the explicit expressions of low-order cumulants and binomial moments. Compared with models under burst approximation, the first-order binomial moment remains exact, whereas the second-order binomial moment generally differs. An upper bound is given for the difference between the second-order binomial moments obtained from two models, mainly using techniques from functional analysis. Additionally, we present the analytical expression of the multi-variable generating function by solving a partial differential equation system.
@@
The article is structured as follows. In subsection 2.1, a general stochastic gene expression model is introduced, in which the gene has arbitrarily many states. In subsection 2.2, the chemical master equation describing the dynamics is presented, and then converted to a partial differential equation system using standard generating function method. Using high-dimensional binomial moment method, an ordinary differential equation hierarchy governing the evolution of binomial moments of mRNA and protein copy number is derived. We present the concise recurrence relations in steady state and design an algorithm for numerical computation. These results are presented in subsection 3.1. In subsection 3.2, we provide an equality to reconstruct the probability mass function from binomial moments. Analytical expressions for low-order cumulants of the number of mRNA and protein molecules are given in subsection 3.3, and are compared with those from models under burst approximation in section 4.
@@ 2.1 Model Description
We now introduce a general stochastic model of gene expression, formulated as a chemical reaction system schemed in (1). In this model, the gene we consider is assumed to have $N$ different states, namely, $\mathcal{S}_{i}\;(1\leq i\leq N)$, and the gene transits arbitrarily among these states in a Markovian manner. In each state, mRNA molecules are transcribed in a different rate and gene-state is allowed to switch upon transcription. Transcribed mRNA molecules are confronted with competing reaction pathways of hydrolysis and translation. Hydrolysis of mRNA molecules occurs at constant rate, independent of the state of gene or number of protein molecules. Specific type of protein molecules are translated from living mRNA molecules at constant rate, and also undergo hydrolysis once produced. Based on the general theory of stochastic chemical reaction kinetics [AppliedStochasticAnalysis2019], (1) defines a continuous-time Markov chain $(S(t),M_{1}(t),M_{2}(t))_{t\geq 0}$ with state space $\{1,2,\cdots,N\}\times\mathbb{N}\times\mathbb{N}$, where $S(t)$, $M_{1}(t)$ and $M_{2}(t)$ denote the state of gene, the number of mRNA molecules in the system, and the number of protein molecules in the system.
@@
$\rightarrowfill@\displaystyle{}\mathrm{{Protein}}{}\mathrel{\mathop{\makebox[0.0pt]{\rightarrowfill@\displaystyle}}\limits^{\mkern 5.0mu{}\mathrm{{}}\mkern 9.0mu}}{}{}\mathrm{\text{$\emptyset$}}$
@@
Detailed stochastic gene expression models tracking the dynamics of both mRNA and protein counts are rarely studied before, because they have long been assumed analytically intractable even in the simplest case. The simplest special case in (1), namely the case where $N=1$, is first analyzed in [Analyticaldistributionsstochastic2008Proc.Natl.Acad.Sci.U.S.A.]. Using generating function method and the method of characteristics for solving linear partial differential equations, the authors of [Analyticaldistributionsstochastic2008Proc.Natl.Acad.Sci.U.S.A.] obtain analytical expression of the generating function but find that it involves an integral that cannot be calculated, namely, $\int\frac{\mathrm{e}^{x}}{x}$. Therefore, they stop further analyzing this model and propose an approximation technique, the burst approximation, that is now predominantly used by the community. As a result, the detailed stochastic gene expression model (1) has long been assumed intractable and has never been thoroughly studied even in any simple cases.
@@ 2.2 Chemical Master Equation
Assume there are no mRNA or protein molecules in the system at time 0 and the gene is in a given state, say, $S_{i}\;(1\leq i\leq N)$. Denote by $\mathbb{P}_{i,j}(m,n;t)\;(1\leq j\leq N,m\in\mathbb{N},n\in\mathbb{N},t\geq 0)$ the probability that at time $t$ there are $m$ mRNA molecules, $n$ protein molecules and the gene is in state $S_{j}$. Define $\mathbb{P}(m,n;t)$ to be a $N\times N$ matrix whose $(i,j)$-th element is $\mathbb{P}_{i,j}(m,n;t)$. The chemical master equation of (1) is
@@
Using standard generating function method, or equivalently, Laplace transform, we define the matrix-form generating function of $\mathbb{P}(m,n,t)$ as
@@
which is exactly the partial differential equation system appeared in our previous work [lu2025stochastickineticsmrnamolecules]. This can be intuitively understood since the fluctuations of protein copy number does not contribute to the fluctuation at mRNA level, while changes in mRNA copy number do affect fluctuations at protein level. This phenomena is widely-known as dynamic disorder [Singlemoleculeapproachdispersed2002J.Chem.Phys.]. Additionally, the probability distribution of protein copy number cannot be treated marginally; instead, the joint probability distribution needs to be considered.
Actually, we also construct the analytical solution to the partial differential equation system (5) using the method of characteristics, which can be found in the Supplementary Material. We note that further analysis based on the analytical solution requires additional development and is not pursued in this article.
@@ 3.1 Two-dimensional Binomial Moment Method
In this and the following section, we aim to establish an efficient numerical scheme for computing the steady-state joint distribution of mRNA and protein copy number. Our approach builds on standard binomial moment method [Binomialmomentequations2011Phys.Rev.Lett., Stochasticanalysiscomplex2012Phys.Rev.Ea, momentconvergencemethodstochastic2016J.Chem.Phys.a], which analyzes the moments of a probability distribution and subsequently reconstructs the corresponding probability mass function. We first obtain the binomial moments in this section.
@@
$\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}m^{p}n^{q}\mathbb{P}(m,n;t)$ [CourseProbabilityTheory2000], therefore binomial moments exist if and only if moments of arbitrary orders exist for each element in $\mathbb{P}(m,n;t)$. The name “binomial” comes from combinatorial coefficients in (6). Additionally, denote the stationary binomial moments by $\mathcal{B}_{p,q}:=\lim_{t\rightarrow\infty}\mathcal{B}_{p,q}(t)$ and the stationary probability mass function by $\mathbb{P}(m,n):=\lim_{t\rightarrow\infty}\mathbb{P}(m,n;t)$.
@@
Note that $\mathcal{B}_{0,0}=\bm{I}_{N}$ by definition and $\mathcal{B}_{p,q}$ is taken as $\bm{0}_{N\times N}$ if $p<0$ or $q<0$. See Supplementary Material for the detailed derivation.
@@
Recall that $D$ is the generator of the Markov chain $(S(t))_{t\geq 0}$. Thereby, $\gamma\bm{I}_{N}-D$ is a strictly diagonally dominant matrix for any $\gamma>0$, and is nonsingular by Lévy-Desplanques theorem [MatrixAnalysis2012]. Additionally, strict diagonal dominance guarantees the numerical stability while performing a LU factorization [MatrixComputations2013]. Thus, the numerical scheme Algorithm S.1 yields binomial moments with high accuracy.
Note that analytical expressions of the binomial moments of mRNA copy number, namely, $\mathcal{B}_{p,0}\;(p\in\mathbb{N})$, agrees with existing results [Multimodalityflexibilitystochastic2013Bull.Math.Biol., lu2025stochastickineticsmrnamolecules, Promotermediatedtranscriptionaldynamics2014Biophys.J.]. However, the derivation here is primarily formal in comparison with our earlier work [lu2025stochastickineticsmrnamolecules], in which the time-dependent binomial moments are derived based on the analytical solution to the chemical master equation, and the stationary expressions are obtained by taking temporal limit.

### v3

# Approximation Error of the Burst Approximation for a Stochastic Gene Expression Model

@@ Abstract
Stochastic modeling of gene expression is a classic problem in theoretical biophysics, and the burst approximation is widely used to simplify gene expression models formulated via the chemical master equation. However, the approximation error has been investigated only for the simplest case. This article proposes and analyzes a general stochastic gene expression model with an arbitrary number of gene states, and quantifies the error introduced by the burst approximation. Using the standard binomial moment method, we derive recurrence relations for binomial moments in steady state. We develop an algorithm to numerically compute binomial moments in a hierarchical manner. In particular, explicit expressions for low-order moments are presented. Compared with surrogate models under the burst approximation, we conclude that the first-order moment of protein counts is preserved, whereas discrepancies generally arise in higher-order moments. By estimating the difference between two second-order moments using functional analysis, we evaluate the validity of the burst approximation.
@@ 1 Introduction
Gene expression is one of the most fundamental processes in molecular biology, through which genetic information is converted into functional gene products. In most cases, gene expression involves two main stages: transcription and translation. During transcription, messenger RNA (mRNA) is synthesized from specific segments of DNA, whereas translation refers to the production of proteins using mRNA as a template. To quantitatively understand and predict this process, it is essential to propose appropriate mathematical models. Notably, intrinsic stochasticity is the key feature of gene expression. Such inherent stochasticity originates from random collisions among molecules, and becomes relatively significant due to low numbers of mRNA and protein molecules in most cells [20]. Randomized expression of genes can be beneficial for biological adaptation to changing environments [30]. However, how these fluctuations propagate within and across cells, and how they are ultimately regulated, remains an open question at the theoretical level.
@@
The primary variables are the mRNA and protein copy numbers during gene expression, whose dynamics may be abstracted as a stochastic process. Note that deterministic approaches like reaction rate equation [8] usually fail to describe this process because of pronounced stochasticity. More specifically, this stochastic process should be a continuous-time process with a discrete state space. The state space should be discrete because the mRNA and protein counts are nonnegative integers. The simplest class of stochastic processes satisfying both conditions consists of continuous-time Markov chains, whose Kolmogorov forward equation is commonly termed the chemical master equation [8, 37] in chemical physics literature. In general, the chemical master equation is an infinite-dimensional coupled system of ordinary differential equations, and can be equivalently converted into a finite-dimensional partial differential equation system using the generating function method or the Laplace transform. When analyzing a dynamical system described by the chemical master equation, neither a universal treatment nor an effective general theory exists, and one usually devises techniques tailored to different problems. In particular, concise analytical results are rarely anticipated, and in many cases, different kinds of approximation techniques are applied [31]. The most widely used approximation method is the diffusion approximation [9], which approximates the chemical master equation with a Markov-type stochastic differential equation [19], conventionally termed the chemical Langevin equation. This procedure approximates a jump process with a diffusion process. Therefore, the diffusion approximation may become inappropriate when the numbers of molecules in the system are extremely low, as the discreteness of the state space becomes significant. Despite the challenges, gene expression models based on the chemical master equation are worth in-depth analysis given their biological importance.
@@
Proposing stochastic gene expression models is straightforward. The simplest two-stage model is analyzed in [33]. In the two-stage model, mRNA molecules are continuously produced according to a Poisson process, and translation, hydrolysis of mRNA, and hydrolysis of protein all occur as memoryless single-step reactions. Using the generating function method and the method of characteristics for solving linear partial differential equations, the authors of [33] obtain a formal analytical expression of the generating function. However, the expression of the generating function contains an incomplete gamma function that cannot be reduced, and an analytical expression for the protein copy number distribution is infeasible [33]. For a long time, results on stochastic gene expression models are largely confined to low-order moments of several toy models [27]. In [15], the binomial moment method [1, 2, 38] is first applied to a full gene expression model with two gene states. Both recurrence relations for binomial moments and analytical expressions for low-order moments are provided [15]. Similar methodology is also used to study a full gene expression model with multiple active and inactive gene states [29]. Related work also includes [21].
@@
Meanwhile, an effective approximation technique, namely, the burst approximation [27, 33], is also available. The burst approximation builds upon the experimental conclusion that mRNAs decay substantially faster than proteins in most cells [33]. Under the burst approximation, the original gene expression models can be greatly simplified and analytical results become readily available [6, 17, 33]. Beyond their roles as approximate substitutes, stochastic gene expression models with translational bursting have independent theoretical value. Particularly, models with multiple gene states under the burst approximation have been studied using the binomial moment method [6], and non-Markovian models can also be studied using techniques from queueing theory [18]. According to the equivalence between queueing systems and gene expression models established in [35], models under the burst approximation can be interpreted as queueing systems with batch arrivals. This perspective may explain the substantial analytical simplifications introduced by the burst approximation, and clarify the theoretical significance of related research. Nevertheless, whether the errors introduced by the burst approximation are controllable has not received much attention. The validity of the burst approximation has only been evaluated in the simplest two-stage model [33], and validation for general models is the main motivation in this article.
@@
Additionally, general numerical methods exist for reaction systems described by the chemical master equation. The stochastic simulation algorithm (SSA) [11] is a numerical method for generating the sample paths of a given continuous-time Markov chain, one realization at a time. Multiple variants of the SSA are also available [10]. In this article, we employ the Python package GillesPy2 [25] to perform the SSA. The finite state projection algorithm (FSP) [26] is also a widely used numerical method, which truncates the chemical master equation and reduces the problem to numerically solving an ordinary differential equation system with constant coefficients. In general, numerical methods including the SSA and the FSP are computationally expensive and quickly become infeasible for large systems. Moreover, the aforementioned numerical methods cannot be used to derive further theoretical results. As a result, theoretically analyzing complete stochastic gene expression models is still important.
@@
In this article, we analyze a general complete gene expression model with multiple gene states, and estimate the error introduced by the burst approximation. Firstly, similar to [15, 29], we apply the high-dimensional binomial moment method to derive recurrence relations for binomial moments of mRNA and protein copy number, thereby yielding exact analytical expressions for binomial moments of arbitrary order. This part is a straightforward generalization of the results in [15, 29]. Next, we design an algorithm that numerically computes, with high accuracy, all binomial moments in a hierarchical way. When binomial moments are obtained, the joint probability mass function of mRNA and protein copy numbers can be constructed directly. In particular, we present explicit expressions for low-order moments. Compared with models under the burst approximation, we conclude that the first-order binomial moment remains exact, whereas the second-order binomial moment generally differs. An upper bound is given for the difference between the second-order binomial moments obtained from two models, mainly using techniques from functional analysis. Finally, we present the analytical expression of the multi-variable generating function by solving a partial differential equation system.
@@
The article is structured as follows. In subsection 2.1, a general stochastic gene expression model is introduced, in which the gene has arbitrarily many states and can transition freely among them. In subsection 2.2, the chemical master equation describing the dynamics is presented, and then converted to a partial differential equation system using the standard generating function method. Using high-dimensional binomial moment method, an ordinary differential equation hierarchy governing the evolution of binomial moments of mRNA and protein copy number is derived. We present the concise recurrence relations in steady state and design an algorithm for numerical computation. These results are presented in subsection 3.1. In subsection 3.2, we provide an equality to reconstruct the probability mass function from binomial moments. Analytical expressions for low-order cumulants of the number of mRNA and protein molecules are given in subsection 3.3, and are compared with those from models under burst approximation in section 4.
@@ 2.1 Model Description
We now introduce a general stochastic model of gene expression, formulated as a chemical reaction system schemed in (1). In this model, the gene we consider is assumed to have $N$ different states, namely, $\mathcal{S}_{i}\;(1\leq i\leq N)$, and the gene transitions arbitrarily among these states in a Markovian manner. In each state, mRNA molecules are transcribed at a different rate and the gene state is allowed to switch upon transcription. Transcribed mRNA molecules are confronted with competing reaction pathways of hydrolysis and translation. Hydrolysis of mRNA molecules occurs at a constant rate, independent of the state of the gene or the number of protein molecules. A specific type of protein molecules is translated from living mRNA molecules at constant rate, and also undergo hydrolysis once produced. Based on the general theory of stochastic chemical reaction kinetics [8], (1) defines a continuous-time Markov chain $(S(t),M_{1}(t),M_{2}(t))_{t\geq 0}$ with state space $\{1,2,\cdots,N\}\times\mathbb{N}\times\mathbb{N}$, where $S(t)$, $M_{1}(t)$ and $M_{2}(t)$ denote the state of the gene, the number of mRNA molecules in the system, and the number of protein molecules in the system.
@@
$\rightarrowfill@\displaystyle{}\mathrm{{Protein}}{}\mathrel{\mathop{\makebox[20.00003pt]{\rightarrowfill@\displaystyle}}\limits^{\mkern 5.0mu{}\mathrm{{}}\mkern 9.0mu}}{}{}\mathrm{\text{$\emptyset$}}$
@@ 2.2 Chemical Master Equation
Assume that there are no mRNA or protein molecules in the system at time 0 and that the gene is in a given state, say, $S_{i}\;(1\leq i\leq N)$. Denote by $\mathbb{P}_{i,j}(m,n;t)\;(1\leq j\leq N,m\in\mathbb{N},n\in\mathbb{N},t\geq 0)$ the probability that at time $t$ there are $m$ mRNA molecules, $n$ protein molecules and the gene is in state $S_{j}$. Define $\mathbb{P}(m,n;t)$ to be a $N\times N$ matrix whose $(i,j)$-th element is $\mathbb{P}_{i,j}(m,n;t)$. The chemical master equation of (1) is
@@
Using the generating function method, or equivalently, the Laplace transform, we define the matrix-form generating function of $\mathbb{P}(m,n,t)$ as
@@
which is exactly the partial differential equation system that appeared in our previous work [22]. This can be intuitively understood since the fluctuations of the protein copy number do not contribute to the fluctuation at mRNA level, while changes in mRNA copy number do affect fluctuations at protein level. This phenomenon is widely known as dynamic disorder [34]. Additionally, the probability distribution of the protein copy number cannot be treated marginally; instead, the joint probability distribution needs to be considered.
Actually, we also construct the analytical solution to the partial differential equation system (5) using the method of characteristics, which can be found in the Supplementary Material. We note that further analysis based on this analytical solution requires additional work and is not pursued in this article.
@@ 3.1 Two-dimensional Binomial Moment Method
In this and the following section, we aim to establish an efficient numerical scheme for computing the steady-state joint distribution of mRNA and protein copy number. Our approach builds on the standard binomial moment method [1, 2, 38], which analyzes the moments of a probability distribution and subsequently reconstructs the corresponding probability mass function. In fact, the binomial moment method can be used to study any first-order reaction system. We first obtain the binomial moments in this section.
@@
$\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}m^{p}n^{q}\mathbb{P}(m,n;t)$ [7], therefore binomial moments exist if and only if moments of arbitrary order exist for each element in $\mathbb{P}(m,n;t)$. The name “binomial” comes from combinatorial coefficients in (6). Additionally, denote the stationary binomial moments by $\mathcal{B}_{p,q}:=\lim_{t\rightarrow\infty}\mathcal{B}_{p,q}(t)$ and the stationary matrix-valued probability mass function by $\mathbb{P}(m,n):=\lim_{t\rightarrow\infty}\mathbb{P}(m,n;t)$.
@@
Note that $\mathcal{B}_{0,0}=\bm{I}_{N}$ by definition and $\mathcal{B}_{p,q}$ is taken as $\bm{0}_{N\times N}$ if $p<0$ or $q<0$. See Supplementary Material for the detailed derivation. (8) agrees with the results in [15, 29].
@@
Recall that $D$ is the generator of the Markov chain $(S(t))_{t\geq 0}$. Therefore, $\gamma\bm{I}_{N}-D$ is a strictly diagonally dominant matrix for any $\gamma>0$, and is nonsingular by Lévy-Desplanques theorem [14]. Additionally, strict diagonal dominance guarantees the numerical stability while performing a LU factorization [12]. Thus, the numerical scheme Algorithm S.1 yields binomial moments with high accuracy.
Note that analytical expressions of the binomial moments of mRNA copy number, namely, $\mathcal{B}_{p,0}\;(p\in\mathbb{N})$, agree with existing results [16, 22, 39]. However, the derivation here is primarily formal in comparison with our earlier work [22], in which the time-dependent binomial moments are derived based on the analytical solution to the chemical master equation, and the stationary expressions are obtained by taking temporal limit.